OpenStudy (anonymous):

Rewrite using rational exponents. \[\sqrt[5]{x^2}\]

OpenStudy (mertsj):

remember: in a rational exponent, the top number is the power and the bottom number is the root.

hero (hero):

Hint: \[\sqrt[a]{x^b} = x^\frac{b}{a}\]

OpenStudy (anonymous):

would the answer be x 2/5?

hero (hero):

x^(2/5) would be the way to write it linear form

OpenStudy (anonymous):

thank you! so how would I rewrite \[(a^\frac{ 3 }{ 4 })^5\] using radicals?

OpenStudy (anonymous):

would it be a^15/4 ?

OpenStudy (anonymous):

@Karina1 ask your question in the little "ask here" box on the left side of the screen

OpenStudy (anonymous):

oh ok

OpenStudy (pixiedust1):

Hero, can you help me when your done?

hero (hero):

@ali1029, \[x^{\frac{a}{b}}= \sqrt[b]{x^a}\]

OpenStudy (anonymous):

how exactly would i change (a^3/4)^5 into the form x^a/b ?

OpenStudy (anonymous):

your answer was correct

OpenStudy (anonymous):

a^15/4 is correct? :D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you both so much!!!! :)

OpenStudy (anonymous):

yw from both of us

hero (hero):

It said to write it in radical form